OpenStudy (anonymous):

The Width Of A Singles Tennis Ball Court Is 27ft . This Is 1Ft More Than 1/3 Of The Length . What Is The Lenght Of The Court ? .

OpenStudy (asnaseer):

call the length of the court L. then you have:\[ 27=1+\frac{L}{3}\] solve to find L.

OpenStudy (anonymous):

very nice

OpenStudy (anonymous):

thus L = 26 *3 or L= 78